# The Multiband Imaging Survey for High-Alpha PlanetS (MISHAPS) I: Preliminary Constraints on the Occurrence Rate of Hot Jupiters in 47 Tucanae

### v1

@@ 5.2 Transit Search
To set the detection threshold, we run the search algorithm on the same sample of $\sim$2,000 injections used to test the models. We plot the differential and cumulative distributions of S/N values of the best boxcar transit model for nights that do contain injected transits, and for nights that do not, in Figure 7. We then find the S/N threshold that allows us to detect a reasonable number of transits, while excluding as many false positives as possible, opting ultimately for $S/N\geq 7$. This choice passes $\sim$47of injected transits, while rejecting $\sim$99of nights without a transit injection.
Figure 7: Histogram showing the S/N of nights with (yellow) and without (blue) injected transits. The dotted red line indicates our chosen S/N7absent7$\geq$ 7≥ 7 threshold. The solid and dashed black lines represent the probability of a lightcurve containing a detection of a given S/N for nights with and without injected transits, respectively.
@@ 5.4 Detection Efficiency
We compute these efficiencies in bins of $R_{\rm p}$, $P$, and $r-$magnitude after applying the color cut, using parameter ranges that encompass those considered by G00, W05, and MW17 ($0.5R_{\rm Jup}\leq R_{\rm p}\leq 2.0R_{\rm Jup}$, 0.5 days $\leq P_{\rm orb}\leq$ 10.0 days). Figure 9 shows the efficiencies as a function of $R_{\rm p}$, $P$, $r-$magnitude, and impact parameter $b$, averaged over the other parameters. Figures 10 and 11 show combined efficiencies over 2D grids of $R_{\rm p}$ vs. $P$ and $R_{\rm p}$ vs. $r-$magnitude for $\epsilon_{\rm det}$ and $\epsilon_{\rm Zoo}$. For the 2D efficiency plots, errors were computed using binomial confidence intervals at a confidence level of 68%, or 1-$\sigma$.
@@ 6 Detailed Candidate Vetting
    Used VARTOOLS’ BLS, AOV, and Lomb-Scargle period searches (Kovács et al., 2002; Schwarzenberg-Czerny, 1989; Devor, 2005; Press et al., 1992; Zechmeister & Kürster, 2009) to obtain preliminary estimates of the target’s period, then folded the lightcurves on those periods to see if a strong transit, EB, or other variable signal emerged.
@@
Of the 39 candidates that made it through the Zooniverse vetting prior to the color cut, only 18 survived this round of analysis. The majority of those rejected — 14 out of 21 — were found to have messy difference images near the target, often with a dipole feature, with many of the dipoles belonging to a very close star within ${\sim}1$ PSF FWHM of the target (i.e., blended in the reference image). An example of a dipole is shown in Figure 12. The dipoles are caused by a quirk of the growth of DECam’s CCD chips. Image artefacts referred to as “tree rings” — visible in flat fields and bright sky backgrounds — are caused by photoelectric imperfections in the DECam CCDs, and result in the effective area of each pixel varying (Plazas et al., 2014; Bernstein et al., 2017, 2018). When the science images are flat-fielded, the tree rings introduce small astrometric shifts in stars positioned on the rings (Baumer et al., 2017). The image alignment step of our difference imaging only accounts for large-scale astrometric distortions, and thus cannot account for these localized shifts on the scale of the entire chip. This shift results in dipoles in our difference imaging because the star may be positioned differently in the reference image than in the individual images relative to an assumed regular pixel grid. Visually, these appear as a positive and negative area in the image where the misaligned star is over-subtracted on one side and under-subtracted on the other. As the aperture is centered on the reference image star, a small shift in the target image can result in some of the positive dipole landing outside the aperture and more of the negative dipole landing inside, which — if this happens in the same sense for a sequence of images — can result in a transit-like systematic in the lightcurve. The reverse case causing an apparent brightening can also look like a partial transit. The chance of such systematics is greatly increased when there is a nearby neighbor with a dipole that might only have the positive or negative side landing in the aperture. We did not find any blended eclipsing binaries through the stacked difference images.
@@
As an extra check on the photometry, we also performed aperture photometry on the cutout images using AstroImageJ (Collins et al., 2017). We compare the original lightcurves, cutout lightcurves, and aperture photometry lightcurves to see whether the eclipse features seen in the lightcurves are consistent or if they disappear, indicating false positives from, e.g., poor initial difference imaging and photometry, or systematic effects caused by the analysis. Finally, we make new in- and out-of-transit image stacks using the target-centered difference images.
@@
N   OTE—For full transits, the $t_{\rm c}$ returned by the transit search algorithm is given. For partial transits,
@@ 6.1 Case Study: MISHAPS_F47T_S11_01004730
The target-centered difference imaging lightcurve and the AstroImageJ aperture photometry lightcurve show a similar event, which lends some credence to the signal being real. However, notably, the $r$-band target-centered lightcurve transit is significantly shallower than the detrended lightcurve’s transit (shown in the middle panel of Figure 2.2). We fit both the original pipeline lightcurve and the target-centered difference image lightcurve with transiting planet models calculated using Batman (Kreidberg, 2015) fits to yield best fit transits with depths of $-12.3$ mmag and $-8.3$ mmag, respectively. As the target-centered photometric extraction uses a larger aperture, it is possible that the inclusion of flux from another star entering the aperture on the reference image could cause the depth to be shallower, but we rule that out as a possibility — the target-centered lightcurve is only ${\sim}0.01$ mag brighter, whereas to cause enough dilution to decrease the transit depth by ${\sim}33$% would require a neighboring star to increase the flux in the aperture by ${\sim}0.31$ mag. This depth difference between the two lightcurves measured with the same data but slightly different techniques was the first piece of evidence to reject the candidate, but we did not deem it conclusive enough on its own. We therefore inspected the individual difference images in a timeseries, as shown in the thin lower panel of Figure 2.2, and found that the variations of strength of the dipole are correlated with the flux, with the dipole nearly disappearing at the bottom of the apparent transit.
In addition to the flagged transit, the lightcurve of this object shows variability at a comparable level on some nights, with some of this variability having the appearance of ingresses, egresses or full transits. Boxcar depths in the detrended lightcurves range from 3 to 9 mmag, with one of these having a boxcar $S/N=6.3$ and the rest lower. Lomb-Scargle, box least squared, and analysis of variance periodograms (Hartman & Bakos, 2016; Zechmeister & Kürster, 2009; Press et al., 1992; Kovács et al., 2002; Schwarzenberg-Czerny, 1989; Devor, 2005) do not reveal any clear periods, which, if the single transit-like signals were periodic and detectable by eye, we would expect to phase together into a clear folded signal.
@@ 6.2 Eclipsing Binaries
Our search and 2-step vetting leave us with 4 EB candidates. We cross-referenced the EB candidates with Weldrake et al. (2004) and to Vizier catalogs containing the targets. One target — MISHAPS_F47T_S5_01002002 — is contained in the OGLE catalog of EBs (Pawlak et al., 2016) and is also flagged as an EB in Gaia DR3 (Babusiaux et al., 2023), but we find no existing literature matches for the other 3. Other catalogs searched included Roman (1987), the Guide Star Catalog (Lasker et al., 2008), Kaluzny et al. (2013), ATLAS-REFCAT 2 (Tonry et al., 2018), Stetson et al. (2019), and Jiménez-Arranz et al. (2023). All four targets lie on 47 Tuc’s binary sequence and show clear indications of eclipses both in their lightcurves and in their stacked in-“transit” difference images. These candidates are all rejected by our color cuts that exclude stars on the binary sequence, but demonstrate the value of MISHAPS’ high-cadence color observations. The flagged portions of their lightcurves, stacked difference images, and reference images are shown in Figure 17
@@ 7 Occurrence Rate Limit
Our goal is to constrain the occurrence rate of hot Jupiters in 47 Tuc. As we have not found any convincing candidates we can only place an upper limit on their occurrence. We estimate our 95% upper limit on the occurrence rate by assuming the number of candidate planets we expect to detect, $N_{\rm exp}$, is Poisson distributed
@@
where $N_{\rm cand}$ is the number of surviving candidates. Setting the probability of finding no planets equal to $(1-0.95)$ and solving for $N_{\rm exp}$ when $N_{\rm cand}=0$, we get $N_{\rm exp}=3$. The number of candidates we expect to detect is set by the number of stars searched $N_{\star}$, the hot Jupiter occurrence rate $f_{\rm HJ}$, and the average total detection efficiency
@@
For these reasons we opt to present our main occurrence rate limit in nearly the full range over which we calculate detection efficiencies, taking the full range of periods $100.5\leq P<$ 10 days, and only a slightly restricted radius range of $0.75\leq R_{\rm p}<2.0~{}R_{\rm Jup}$. We decide to ignore the 0.50.50.50.50.750.750.750.75 $R_{\rm Jup}$ range of planet radii because our average detection efficiency drops precipitously by a factor of ${\sim}5.5$ relative to the 0.750.750.750.751.01.01.01.0 $R_{\rm Jup}$ bin, and because we expect very few hot Jupiters to have radii below $0.75$ $R_{\rm Jup}$ if their incident flux is above ${\sim}2\times 10^{8}$ erg s-1 cm-2 (e.g., Demory & Seager, 2011; Thorngren & Fortney, 2018). For a planet in a 10 day orbit around an $r=21$ star in 47 Tuc, we compute a minimum incident flux of $2.5\times 10^{8}$ erg s-1 cm-2 using the MIST version 1.2 isochrone with [Fe/H]=-0.75 dex at a distance of 4.45 kpc and with no extinction (Dotter, 2016; Choi et al., 2016).
@@
Using our detection efficiencies for the 19,930 stars passing our target selection cuts, we estimate $N_{1}=693$ for $0.75\leq R_{\rm p}<2.0~{}R_{\rm Jup}$ and $100.5\leq P<$ 10 days, which yields the upper limit on the occurrence rate of hot Jupiters in 47 Tuc
@@
[…]
@@
It is also useful to compare our limits to the occurrence rate in the solar neighborhood. Fressin et al. (2013) present occurrence rates using data from the Kepler spacecraft corrected for false positives. In the period ranges $0.8$––$3.4$ days and 0.80.80.80.810101010  days, for planets with radius $0.53\leq R_{\rm p}<1.96R_{\text{Jup}}$, they measure hot Jupiter occurrence rates of $0.082\pm 0.019$% and $0.43\pm 0.05$%. Over these period and $R_{\rm p}$ ranges we compute $N_{1}=1178$ and $515$, and place limits of $f_{\rm HJ}<0.25\%$ and $<0.58\%$, respectively.
@@
To enable more detailed comparisons with future work, we provide with this article the results of our injection-recovery tests for all injections at the NOIRLab Astro Data Lab. For calculations that do not require data processing, we also present our occurrence rate limits for bins of planet radius and period in Figure 18. To combine cells in this plot, the individual limits can be converted to $N_{1}=3/f_{\rm HJ,max}$, averaged, and then converted back to a limit with equation 21.
@@ 47 Tuc estimates
This work 0.750.750.750.752.02.02.02.0 0.50.50.50.510.010.010.010.0 $19,930$ $692.8$ $<0.43$%
Gilliland et al. (2000) 0.80.80.80.82.02.02.02.0 0.50.50.50.58.38.38.38.3 $34,091$ $1888.9$ $<0.16$%
MW17 Kepler hot Jupiters $34,091$ $1222.2$ $<0.25$%
This work 0.80.80.80.82.02.02.02.0 0.50.50.50.58.38.38.38.3 $19,930$ $830.1$ $<0.36$%
This work + G00 0.80.80.80.82.02.02.02.0 0.50.50.50.58.38.38.38.3 $54,021$ $2719.0$ $<0.11$%
Weldrake et al. (2005) $1.3$ 1.01.01.01.010.010.010.010.0 $21,920$ $377.7$ $<0.79$%
This work 1.21.21.21.21.41.41.41.4 1.01.01.01.010.010.010.010.0 $19,930$ $564.6$ $<0.53$%
This work 0.530.530.530.531.961.961.961.96 0.80.80.80.83.43.43.43.4 $19,930$ $1177.8$ $<0.25$%
This work 0.530.530.530.531.961.961.961.96 0.80.80.80.810.010.010.010.0 $19,930$ $514.6$ $<0.58$%
@@ 47 Tuc predictions
Masuda & Winn (2017) $0.880.57\leq M_{\star}<$ 0.88 0.80.80.80.82.02.02.02.0 0.50.50.50.58.38.38.38.3 $0.18$%
Masuda & Winn (2017) $0.880.57\leq M_{\star}<$ 0.88 0.80.80.80.82.02.02.02.0 0.680.680.680.6810.010.010.010.0 $0.24_{-0.09}^{+0.10}$%
Fressin et al. (2013) $0.880.57\leq M_{\star}<$ 0.88 & [Fe/H]$=-0.78$ 0.530.530.530.531.961.961.961.96 0.80.80.80.810.010.010.010.0 $0.028$%
Fressin et al. (2013) $0.880.57\leq M_{\star}<$ 0.88 & [$\alpha$/H]$=-0.48$ 0.530.530.530.531.961.961.961.96 0.80.80.80.810.010.010.010.0 $0.055$%
Note. — Occurrence rates for this work are computed for different ranges of planet radius and period to match other studies. The 47 Tuc predictions list two different estimates from MW17 that are their estimates of the Kepler-field occurrence rates for masses matching 47 Tuc’s main sequence stars searched by G00 but with different period ranges, and assume no metallicity dependence. The , predictions scale the Kepler occurrence rates measured by , for the same host mass range by the joint mass-metallicity planet occurrence dependence estimated by Johnson et al. (2010) for 47 Tuc’s [Fe/H], and assuming that [$\alpha$/Fe] can be substituted for [Fe/H] in the relation.

### v2

@@ 5.2 Transit Search
To set the detection threshold, we run the search algorithm on the same sample of $\sim$2,000 injections used to test the models. We plot the differential and cumulative distributions of S/N values of the best boxcar transit model for nights that do contain injected transits, and for nights that do not, in Figure 7. We then find the S/N threshold that allows us to detect a reasonable number of transits, while excluding as many false positives as possible, opting ultimately for $S/N\geq 7$. This choice passes $\sim$47% of injected transits, while rejecting $\sim$99% of nights without a transit injection.
[…]
@@ 5.4 Detection Efficiency
We compute these efficiencies in bins of $R_{\rm p}$, $P$, and $r$-magnitude after applying the color cut, using parameter ranges that encompass those considered by G00, W05, F13, and MW17 ($0.5R_{\rm Jup}\leq R_{\rm p}\leq 2.0R_{\rm Jup}$, 0.5 days $\leq P_{\rm orb}\leq$ 10.0 days). Figure 9 shows the efficiencies as a function of $R_{\rm p}$, $P$, $r$-magnitude, and impact parameter $b$, averaged over the other parameters. Figures 10 and 11 show combined efficiencies over 2D grids of $R_{\rm p}$ vs. $P$ and $R_{\rm p}$ vs. $r$-magnitude for $\epsilon_{\rm det}$ and $\epsilon_{\rm Zoo}$. For the 2D efficiency plots, errors were computed using binomial confidence intervals at a confidence level of 68%, or 1-$\sigma$.
@@ 6 Detailed Candidate Vetting
    Used VARTOOLS’ BLS, AOV, and Lomb-Scargle period searches (G. Kovács et al., 2002; A. Schwarzenberg-Czerny, 1989; J. Devor, 2005; W. H. Press et al., 1992; M. Zechmeister & M. Kürster, 2009) to obtain preliminary estimates of the target’s period, then folded the lightcurves on those periods to see if a strong transit, EB, or other variable signal emerged.
@@
Of the 39 candidates that made it through the Zooniverse vetting prior to the color cut, only 18 survived this round of analysis. The majority of those rejected — 14 out of 21 — were found to have messy difference images near the target, often with a dipole feature, with many of the dipoles belonging to a very close star within ${\sim}1$ PSF FWHM of the target (i.e., blended in the reference image). An example of a dipole is shown in Figure 12. The dipoles are caused by a quirk of the growth of DECam’s CCD chips. Image artefacts referred to as “tree rings” — visible in flat fields and bright sky backgrounds — are caused by photoelectric imperfections in the DECam CCDs, and result in the effective area of each pixel varying (A. A. Plazas et al., 2014; G. M. Bernstein et al., 2017, 2018). When the science images are flat-fielded, the tree rings introduce small astrometric shifts in stars positioned on the rings (M. Baumer et al., 2017). The image alignment step of our difference imaging only accounts for large-scale astrometric distortions, and thus cannot account for these localized shifts on the scale of the entire chip. This shift results in dipoles in our difference imaging because the star may be positioned differently in the reference image than in the individual images relative to an assumed regular pixel grid. Visually, these appear as a positive and negative area in the image where the misaligned star is over-subtracted on one side and under-subtracted on the other. As the aperture is centered on the reference image star, a small shift in the target image can result in some of the positive dipole landing outside the aperture and more of the negative dipole landing inside, which — if this happens in the same sense for a sequence of images — can result in a transit-like systematic in the lightcurve. The reverse case causing an apparent brightening can also look like a partial transit. The chance of such systematics is greatly increased when there is a nearby neighbor with a dipole that might only have the positive or negative side landing in the aperture. We did not find any blended eclipsing binaries through the stacked difference images.
@@
As an extra check on the photometry, we also performed aperture photometry on the cutout images using AstroImageJ (K. A. Collins et al., 2017). We compare the original lightcurves, cutout lightcurves, and aperture photometry lightcurves to see whether the eclipse features seen in the lightcurves are consistent or if they disappear, indicating false positives from, e.g., poor initial difference imaging and photometry, or systematic effects caused by the analysis. Finally, we make new in- and out-of-transit image stacks using the target-centered difference images.
@@
a    NOTE—For full transits, the $t_{\rm c}$ returned by the transit search algorithm is given. For partial transits,
@@ 6.1 Case Study: MISHAPS_F47T_S11_01004730
The target-centered difference imaging lightcurve and the AstroImageJ aperture photometry lightcurve show a similar event, which lends some credence to the signal being real. However, notably, the $r$-band target-centered lightcurve transit is significantly shallower than the detrended lightcurve’s transit (shown in the middle panel of Figure 2.2). We fit both the original pipeline lightcurve and the target-centered difference image lightcurve with transiting planet models calculated using Batman (L. Kreidberg, 2015) fits to yield best fit transits with depths of 12.3-12.3 mmag and 8.3-8.3 mmag, respectively. As the target-centered photometric extraction uses a larger aperture, it is possible that the inclusion of flux from another star entering the aperture on the reference image could cause the depth to be shallower, but we rule that out as a possibility — the target-centered lightcurve is only ${\sim}0.01$ mag brighter, whereas to cause enough dilution to decrease the transit depth by ${\sim}33$% would require a neighboring star to increase the flux in the aperture by ${\sim}0.31$ mag. This depth difference between the two lightcurves measured with the same data but slightly different techniques was the first piece of evidence to reject the candidate, but we did not deem it conclusive enough on its own. We therefore inspected the individual difference images in a timeseries, as shown in the thin lower panel of Figure 2.2, and found that the variations of strength of the dipole are correlated with the flux, with the dipole nearly disappearing at the bottom of the apparent transit.
In addition to the flagged transit, the lightcurve of this object shows variability at a comparable level on some nights, with some of this variability having the appearance of ingresses, egresses or full transits. Boxcar depths in the detrended lightcurves range from 3 to 9 mmag, with one of these having a boxcar $S/N=6.3$ and the rest lower. Lomb-Scargle, box least squared, and analysis of variance periodograms (J. D. Hartman & G. Á. Bakos, 2016; M. Zechmeister & M. Kürster, 2009; W. H. Press et al., 1992; G. Kovács et al., 2002; A. Schwarzenberg-Czerny, 1989; J. Devor, 2005) do not reveal any clear periods, which, if the single transit-like signals were periodic and detectable by eye, we would expect to phase together into a clear folded signal.
@@ 6.2 Eclipsing Binaries
Our search and 2-step vetting leave us with 4 EB candidates. We cross-referenced the EB candidates with D. T. F. Weldrake et al. (2004) and to Vizier catalogs containing the targets. One target — MISHAPS_F47T_S5_01002002 — is contained in the OGLE catalog of EBs (M. Pawlak et al., 2016) and is also flagged as an EB in Gaia DR3 (C. Babusiaux et al., 2023), but we find no existing literature matches for the other 3. Other catalogs searched included N. G. Roman (1987), the Guide Star Catalog (B. M. Lasker et al., 2008), J. Kaluzny et al. (2013), ATLAS-REFCAT 2 (J. L. Tonry et al., 2018), P. B. Stetson et al. (2019), and Ó. Jiménez-Arranz et al. (2023). All four targets lie on 47 Tuc’s binary sequence and show clear indications of eclipses both in their lightcurves and in their stacked in-“transit” difference images. These candidates are all rejected by our color cuts that exclude stars on the binary sequence, but demonstrate the value of MISHAPS’ high-cadence color observations. The flagged portions of their lightcurves, stacked difference images, and reference images are shown in Figure 17
@@ 7 Occurrence Rate Limit
Our goal is to constrain the occurrence rate of hot Jupiters in 47 Tuc. For the purpose of this paper, we define our occurrence rate $f_{\rm HJ}$ as the average number of planets per star, assuming that period and radius are distributed uniformly in each parameter across a specified range, i.e.,
$f_{\rm HJ}=\int_{R_{\rm p,min}}^{R_{\rm p,max}}\int_{P_{\rm min}}^{P_{\rm max}}\frac{d^{2}N}{dR_{\rm p}dP}dPdR_{\rm p},$ (18)
where $R_{\rm p}$ is the planet radius and $P$ is the orbital period, and $d^{2}N/dR_{\rm p}dP$ is the joint period-radius distribution.
As we have not found any convincing candidates, we can only place an upper limit on their occurrence. We estimate our 95% upper limit on the occurrence rate by assuming the number of candidate planets we expect to detect, $N_{\rm exp}$, is Poisson distributed
@@
where $N_{\rm cand}$ is the number of surviving candidates. Setting the probability of finding no planets equal to (10.95)(1-0.95) and solving for $N_{\rm exp}$ when $N_{\rm cand}=0$, we get $N_{\rm exp}=3$. The number of candidates we expect to detect is set by $f_{\rm HJ}$, $N_{\star}$, and $\epsilon_{\rm total}$ as
@@
For these reasons we opt to present our main occurrence rate limit in nearly the full range over which we calculate detection efficiencies, taking the full range of periods $0.5\leq P<10$ days, and only a slightly restricted radius range of $0.75\leq R_{\rm p}<2.0~R_{\rm Jup}$. We decide to ignore the 0.50.50.750.75 $R_{\rm Jup}$ range of planet radii because our average detection efficiency drops precipitously by a factor of ${\sim}5.5$ relative to the 0.750.751.01.0 $R_{\rm Jup}$ bin, and because we expect very few hot Jupiters to have radii below $0.75$ $R_{\rm Jup}$ if their incident flux is above ${\sim}2\times 10^{8}$ erg s-1 cm-2 (e.g., B.-O. Demory & S. Seager, 2011; D. P. Thorngren & J. J. Fortney, 2018). For a planet in a 10 day orbit around an $r=21$ star in 47 Tuc, we compute a minimum incident flux of $2.5\times 10^{8}$ erg s-1 cm-2 using the MIST version 1.2 isochrone with [Fe/H]=-0.75 dex at a distance of 4.45 kpc and with no extinction (A. Dotter, 2016; J. Choi et al., 2016).
@@
Using our detection efficiencies for the 19,930 stars passing our target selection cuts, we estimate $N_{1}=693$ for $0.75\leq R_{\rm p}<2.0~R_{\rm Jup}$ and $0.5\leq P<10$ days, which yields the upper limit on the occurrence rate of hot Jupiters in 47 Tuc
@@
G00 used injection-recovery simulations of their Hubble data for planets with $0.8\leq R_{\rm p}<2.0~R_{\rm Jup}$ and $0.5\leq P<8.3$ days to predict their expected yield under the assumption of an occurrence rate $f_{\rm HJ}=0.8$1.01.0%, $N_{\rm exp}=17$. Taking the center of their occurrence rate range, substituting equation 21 into equation 20, and rearranging, we can compute their survey sensitivity
@@
It is also useful to compare our limits to the occurrence rate in the solar neighborhood. F. Fressin et al. (2013) present occurrence rates using data from the Kepler spacecraft corrected for false positives. In the period ranges 0.80.8––$3.4$ days and 0.80.81010  days, for planets with radius $0.53\leq R_{\rm p}<1.96R_{\text{Jup}}$, they measure hot Jupiter occurrence rates of $0.082\pm 0.019$% and $0.43\pm 0.05$%. Over these period and $R_{\rm p}$ ranges we compute $N_{1}=1178$ and $515$, and place limits of $f_{\rm HJ}<0.25\%$ and $0.58\%$, respectively.
@@
To enable more detailed comparisons with future work, we provide with this article the results of our injection-recovery tests for all injections at the NOIRLab Astro Data Lab [^5] https://datalab.noirlab.edu/data/mishaps, and a sample dataset on Zenodo [^6] https://doi.org/10.5281/zenodo.15857086. For calculations that do not require data processing, we also present our occurrence rate limits for bins of planet radius and period in Figure 18. To combine cells in this plot, the individual limits can be converted to $N_{1}=3/f_{\rm HJ,max}$, averaged, and then converted back to a limit with equation 22.
@@ 47 Tuc estimates
This work 0.750.752.02.0 0.50.510.010.0 $19,930$ $692.8$ <0.43<0.43%
R. L. Gilliland et al. (2000) 0.80.82.02.0 0.50.58.38.3 $34,091$ $1888.9$ <0.16<0.16%
MW17 Kepler hot Jupiters $34,091$ $1222.2$ <0.25<0.25%
This work 0.80.82.02.0 0.50.58.38.3 $19,930$ $830.1$ <0.36<0.36%
This work + G00 0.80.82.02.0 0.50.58.38.3 $54,021$ $2719.0$ <0.11<0.11%
D. T. F. Weldrake et al. (2005) $1.3$ 1.01.010.010.0 $21,920$ $377.7$ <0.79<0.79%
This work 1.21.21.41.4 1.01.010.010.0 $19,930$ $564.6$ <0.53<0.53%
This work 0.530.531.961.96 0.80.83.43.4 $19,930$ $1177.8$ <0.25<0.25%
This work 0.530.531.961.96 0.80.810.010.0 $19,930$ $514.6$ <0.58<0.58%
@@ 47 Tuc predictions
K. Masuda & J. N. Winn (2017) $0.57\leq M_{\star}<0.88$ 0.80.82.02.0 0.50.58.38.3 0.180.18%
K. Masuda & J. N. Winn (2017) $0.57\leq M_{\star}<0.88$ 0.80.82.02.0 0.680.6810.010.0 $0.24_{-0.09}^{+0.10}$%
F. Fressin et al. (2013) $0.57\leq M_{\star}<0.88$ & [Fe/H]=0.78=-0.78 0.530.531.961.96 0.80.810.010.0 0.0280.028%
F. Fressin et al. (2013) $0.57\leq M_{\star}<0.88$ & $\alpha$/H]=0.48=-0.48 0.530.531.961.96 0.80.810.010.0 0.0550.055%
Note. — Occurrence rates for this work are computed for different ranges of planet radius and period to match other studies. The 47 Tuc predictions list two different estimates from MW17 that are their estimates of the Kepler-field occurrence rates for masses matching 47 Tuc’s main sequence stars searched by G00 but with different period ranges, and assume no metallicity dependence. The Kepler field predictions scale the occurrence rates measured by F13 for the same host mass range by the joint mass-metallicity planet occurrence dependence estimated by J. A. Johnson et al. (2010) for 47 Tuc’s [Fe/H], and assuming that $\alpha$/H] can be substituted for [Fe/H] in the relation.